# Cooling of an objetct by forced convection

### v1

@@ 2 Theoretical framework
If the radiation contribution is significant, we have to add a term to Eq. 1 and it becomes
@@
In this work we use a thermal camera to measure the temperature in function of time of a vertical plate, while its temperature decrease in a forced convection. We calculated the characteristic cooling time $\tau$ for different air velocities and compare with the previous scaling relation given by Eq. 6.
@@ 4 Results and discussion
Using the thermal camera record we obtained the temperature of the plate as it cooled over time for each air velocity, and calculated $log(T_{o}-T_{e})$ in order to analyze the consistency of the experimental data with the linearization of Eq. 3. Fig. 2 shows this relation, and the linear fit of the experimental values for each cooling. Given that the correlation coefficient is $r=0.997$ for $v=5.5$m/s, and $r=0.999$ for the other velocities, is not necessary to take into account the radiation heat loss. We even compared the two terms on the right hand side of the Eq. 4 for different temperature values, and in all the cases we obtain that $\alpha_{conv}\cdot A\cdot(T_{o}-T_{e})>>\varepsilon\cdot\sigma\cdot A\cdot(T_{o}^{4}-T_{e}^{4})$.
@@
$\tau\frac{\kappa d^{2}}{c}=f\left(\frac{vcd}{\kappa}\right)$ (7)
@@
Figure 3: Log-log plot of the characteristic cooling time, $\tau$, as a function of mean stream velocity. Each characteristic time value was obtained from the slope of of the plot $log(T_{o}-T_{e})=f(t)$. The blue line represents the linear fit of the experimental values, obtained through the least squares method. Furthermore, the shaded blue area represents the 95% confidence interval of this fit. The slope of this graph reveals that the relationship $\tau\propto v^{-1/2}$ is reasonably valid within the margins of uncertainty.
@@ 5 Conclusions
In this work we experimentally study the forced convection of a metal plate placed in front of an air stream. The plate was heated to a predetermined temperature and by means of a thermal camera we studied the characteristic cooling time in relation to the average velocity of the air stream with the help of a thermal camera. The experimental results reveal that $\tau$ for each air velocity and the results were consistent with $\tau\propto v^{-1/2}$.

### v2

@@ 2 Theoretical framework
obtained from the temporal evolution of the temperature linearized as
$\log(T_{o}-T_{e})=\log(T_{i}-T_{e})-\frac{t}{\tau}.$ (4)
It is worth mentioning that in a situation where the radiation contribution is significant we must add a term to Eq. 1 to become
@@
In this work we use a thermal camera to measure the temperature of the vertical plate as a function of time of a vertical plate under the effect of forced convection originated by an air stream. For each air velocity we calculate the characteristic cooling time $\tau$ and compare with the previous scaling relation given by Eq. 7.
@@ 4 Results and discussion
Using the thermal camera record we obtained the temperature of the plate as it cooled over time for each air velocity, and calculated $\log(T_{o}-T_{e})$ in order to analyze the consistency of the experimental data with the linearization of Eq. 3. Fig. 2 shows this relation, and the linear fit of the experimental values for each cooling. Given that the correlation coefficient is $r=0.997$ for $v=5.5$m/s, and $r=0.999$ for the other velocities, is not necessary to take into account the radiation heat loss. We even compared the two terms on the right hand side of the Eq. 5 for different temperature values, and in all the cases we obtain that $\alpha_{conv}\cdot A\cdot(T_{o}-T_{e})>>\varepsilon\cdot\sigma\cdot A\cdot(T_{o}^{4}-T_{e}^{4})$.
@@
$\tau\frac{\kappa}{cd^{2}}=f\left(\frac{vcd}{\kappa}\right)$ (8)
@@
Figure 3: Log-log plot of the characteristic cooling time, $\tau$, as a function of mean stream velocity. The blue line represents the linear fit of the experimental values, obtained through the least squares method. Furthermore, the shaded blue area represents the 95% confidence interval of this fit. The slope of this graph reveals that the relationship $\tau\propto v^{-1/2}$ is reasonably valid within the margins of uncertainty.
@@ 5 Conclusions
In this work we experimentally study the forced convection of a metal plate placed in front of an air stream. The plate was heated to a predetermined temperature and by means of a thermal camera we studied the characteristic cooling time in relation to the average velocity of the air stream. The experimental results reveal that $\tau$ for each air velocity and the results were consistent with $\tau\propto v^{-1/2}$.